A canal system is shown in the figure. Water flows from A to B through two channels. Gates $G_1$ and $G_2$ are operated independently to regulate the flow. Probability of $G_1$ to be open is 10% while that of $G_2$ is 20%. The probability that water will flow from A to B is
To calculate the probability of water flowing from A to B, we need to consider the scenarios in which the gates $G_1$ and $G_2$ are open because water can only flow if at least one gate is open.
Given:
The probability that both gates are closed is calculated as follows:
Therefore, the probability that both gates are closed is the product of both individual probabilities:
\(P(\text{Both Closed}) = 0.90 \times 0.80 = 0.72\)
Thus, the probability that at least one gate is open, allowing water to flow, is:
\(P(\text{At least one open}) = 1 - P(\text{Both Closed}) = 1 - 0.72 = 0.28\)
So, the probability that water will flow from A to B is 28%.
Therefore, the correct answer is 28%.
Consider a Markov chain having state space $S = \{1,2,3,4\}$ with transition probability matrix $P = (p_{i,j})$ given by
$$\begin{matrix} & \begin{matrix} 1 & \quad 2 & \quad 3 & \quad 4 \end{matrix} \\ \begin{matrix} 1 \\ 2 \\ 3 \\ 4 \end{matrix} & \begin{bmatrix} 1/2 & 0 & 1/2 & 0 \\ 1/4 & 1/4 & 1/4 & 1/4 \\ 1/3 & 0 & 1/3 & 1/3 \\ 1/2 & 0 & 1/2 & 0 \end{bmatrix} \end{matrix}$$
Then